# GHP-MOFassemble: Diffusion modeling, high throughput screening, and molecular dynamics for rational discovery of novel metal-organic frameworks for carbon capture at scale

### v1

@@ 3.1 Decomposing MOF linkers into molecular fragments
    Fragment: We fragment the linkers produced in step (ii) to obtain their chemically relevant fragment-connection atom pairs, which we assemble into a molecular fragment dataset.
The extraction of linkers in step (ii) is straightforward because the MOFid provided by hMOF for each MOF specifies the SMILES strings of its constituent inorganic and organic building blocks, including its linkers, as well as a format signature and a topology code [39]. Together, these elements uniquely define the topology and building blocks of a given MOF structure.
Step (iii) use the Matched Molecular Pairs Algorithm (MMPA) [40] as implemented in DeLinker [41] to generate molecular fragments of a given molecule by breaking chemical bonds between atom pairs. We set the minimum number of connection atoms, the minimum fragment size, and the minimum path length to 3, 5, and 2, respectively. Moreover, we only consider the case where the fragments are at least two atoms away from each other. The chemically relevant fragment-connection atoms pairs are then used to form the molecular fragment dataset.
Figure 1: Schematic representation of Decompose, the first component of our GHP-MOFassemble framework. It consists of three steps, which we showcase for Zn tetramer-pcu MOFs. First, the Select step selects the Zn tetramer-pcu MOFs in hMOF that are high-performing. Then the Extract step extracts unique linkers from those MOFs. Finally, the Fragment step generates unique linker fragments. The color scheme of elements is: carbon in grey, oxygen in red, nitrogen in blue, and hydrogen in white.
3.2 Generating new MOF structures
@@
We apply the pre-trained diffusion model DiffLinker [14] to the molecular fragments output by the Decompose component to generate novel MOF linkers. This model connects the molecular fragments supplied as input (also known as context) with a specified number of sampled atoms. We vary the number of sampled atoms from five to 10, a range chosen because we found that with fewer than five sampled atoms, straight chains or branched chains may be obtained, while with more than 10, ring structures may be present. We found that the range of five to 10 sampled atoms resulted in novel and chemically diverse MOF linkers.
@@
DiffLinker applies a denoising process to determine the atomic species and Cartesian coordinates of sampled atoms by using a decoder neural network architecture named E(3)-Equivariant Graph Neural Network (EGNN[42]. This graph neural network predicts zero-mean centered Cartesian coordinates noise and one-hot encoding noise of atomic species, and accounts for equivariance due to translation, rotation, and reflection of molecules, i.e., Euclidean group 3 (E(3)). Since the pre-trained DiffLinker model was trained on the GEOM dataset [43], it enables sampling of connection atoms with high chemical and structural diversity.
@@
DiffLinker outputs the 3D coordinates and the atomic species of heavy atoms in the linker molecules, rather than SMILES strings or 2D graphs. Since hydrogen atoms are implicit in the DiffLinker model, their 3D coordinates are not output by the model. To generate the spatial configurations of all-atom molecules, we employ openbabel to convert the DiffLinker generated molecules to SMILES strings, which in turn are used to generate the 3D configurations of linkers with explicit hydrogen atoms. Openbabel uses distance-based heuristics to determine bond connectivity in a given molecule [44], which is critical to the assignment of the number of hydrogen atoms. After the hydrogen addition step, we identify the dummy atoms which contains information about how the linkers connect with metal nodes.
@@
The diffusion process involves two consecutive steps. The first adds Gaussian noise to the original data (i.e., x), yielding noisy input (i.e., $z_{t}$ at time t). Then, the denoising step applies the neural network-based noise removal operation. Mathematically, the following Markovian properties and equations are satisfied:
@@
3.2.2 Screen and Evaluate.
To ensure consistency of elements in hMOF linkers and generated linkers, we manuallyfilter out linkers that contain elements not present in the hMOF dataset. Since the pre-trained DiffLinker model we used in this work was trained on the GEOM dataset, a total of nine heavy elements may be sampled, including C, N, O, F, P, S, Cl, Br, and I. Among these elements, S, Br, and I do not appear in the hMOF dataset, therefore we remove linkers that contain these three elements. For each molecular fragment, we then perform sampling 20 times. Each sampling step may yield a different set of molecules because of the random nature of the denoising process. The probabilistic nature of the diffusion model allows it to generate linkers from an extensive linker design space beyond that of the hMOF dataset.
@@
The SAscore, as defined by Ertl and Schuffenhauer [47], is based on analysis of one million PubChem molecules, and combines fragment contributions from molecule substructures with a complexity penalty that accounts for molecular size and for structural features of molecules such as the presence of rings. The SCscore is computed by using a neural network trained on 12 million chemical reactions from the Reaxys database to estimate the number of reaction steps required to produce a molecule [46]. For both SAscore and SCscore, the higher the values are, the more difficult it is to synthesize the linker, hence less desirable.
To evaluate the capability of GHP-MOFassemble to generate a valid, novel, and diverse set of linkers, we leverage the MOSES framework [48] to compute three additional metrics: the fraction with valid SMILES strings (validity), the fraction that are unique (uniqueness), and their dissimilarity (internal diversity). We compute the fraction of linkers that are unique by comparing their canonical SMILES strings, and use their average pairwise molecular Tanimoto dissimilarity [49] as measures of diversity. We compute the diversity of the generated linker set by using two internal diversity scores $(\textrm{IntDiv}_{1}(G),\textrm{IntDiv}_{2}(G))$. We present the results of these metrics in Section 4.2.
Figure 2: The Generate component of the GHP-MOFassemble framework involves Diffuse and Denoise, Screen, and Assemble steps, as shown here for Zn tetramer-pcu MOFs. First, we generate new linkers via the Diffuse and Denoise step. We then add hydrogen atoms to ensure correct valency of the generated linkers. Next, we identify dummy atoms by replacing the two carbon atoms in the carboxyl groups of each generated linker with dummy atoms. Linkers with dummy atoms then undergo the Screen step, which removes linkers with S, Br and I. These three elements are present in the GEOM dataset but not in hMOF dataset, and thus we filter out linkers with these three elements. As a result, this step reduces the number of potential linkers to 5,162. The generated linkers’ molecular statistics are quantified using five metrics, including SCscore, SAscore, validity, uniqueness, and internal diversity. Finally, in Assemble step, we build 40,000 new MOFs with the Zn tetramer node in pcu topology. As before, the color code used for the atoms shown is: carbon in grey, oxygen in red, nitrogen in blue, and hydrogen in white.
3.2.3 Assemble.
Once we have generated linkers that meet the five metrics mentioned above, we can assemble them by connecting the node and linker building blocks. To construct MOFs, we need to guide this assembly process. We do that by using dummy atoms, which indicate the points where the building blocks are to be connected. In practice, this is done as follows. Our parsing of MOFids generates, for MOFs with Zn tetramer nodes, linkers that carry two carboxyl groups, which by definition are part of metal nodes instead of linkers. Such wrong assignment of carboxyl groups is due to how the MOFid algorithm parses MOF structures. To reflect the correct molecular structure of linkers, the two carbon atoms in the carboxyl groups are identified as dummy atoms, and the redundant four oxygen atoms and two hydrogen atoms are removed. An illustration of dummy atom identification for linkers containing carboxyl groups is shown in the left panel of Figure 11.
More than half of the hMOF structures are catenated MOFs, i.e., MOFs with interpenetrated lattices. By varying the level of interpenetration, it is possible to generate MOFs with different pore sizes, with a higher catenation level generally corresponding to smaller pores. We denote the four catenation levels in hMOF, with increasing number of interpenetrated lattices, as cat0, cat1, cat2 and cat3. To generate MOFs with high structural diversity, we applied the site translation method as implemented in Pymatgen [50] to generate MOFs with different catenation levels. For catenated MOF structures, we keep the spacing between interpenetrated lattices the same, so as to ensure equal pore sizes.
3.3 Applying regression model to estimate MOF $\textrm{CO}_{2}$ capacity
The Predict component of the GHP-MOFassemble framework uses a regression model to estimate the CO2 capacity of newly generated pcu MOF structures. We use for this purpose a modified version of the Crystal Graph Convolutional Neural Network (CGCNN) model, developed in our previous work [18], which uses an adjacency list to format node and edge embeddings, rather than the adjacency matrix format of the original CGCNN, a change that enhances model training speed, training stability, and prediction accuracy. We trained this modified CGCNN model on all hMOF structures, with a target property of CO2 capacity at 0.1 bar.
We create our ensemble regression model by training the modified CGCNN model three times and averaging the results. We applied the resulting model to the assembled pcu MOF structures produced by the Generate step of Section 3.2 to identify those with a predicted CO2 capacity higher than 2 mmol/g. These MOFs (see Section 4.4.2) are candidates for further investigation by, for example, by performing CO2 capacity validation using Grand Canonical Monte Carlo (GCMC) simulations.
4 Results
By combining the generative power of diffusion models and the computing power of supercomputers, the GHP-MOFassemble framework enables rapid generation and assembly of chemically diverse MOF structures. Using the hMOF dataset as a starting point, our framework accelerates the computational design of high-performing MOF structures for carbon capture by high-throughput screening of novel MOF structures with high CO2 capacities. Our key findings are summarized below.
4.1 Analysis of the hMOF dataset
The three most frequent node-topology pairs in the hMOF dataset, accounting for around 74% of its MOFs, are the Cu paddlewheel-pcu, Zn paddlewheel-pcu, and Zn tetramer-pcu structures shown in the left panel of Figure 3. As summarized in Table 1, Zn tetramer-pcu is the most abundant, followed by Cu paddlewheel-pcu and then Zn paddlewheel-pcu. The cumulative distribution functions of the CO2 capacities of these three types of MOF structure are shown in the right panel of Figure 3. Only the 78,238 MOFs from hMOF with correctly parsed MOFids and valid SMILES were used for analysis. We gather the molecular fragments numbers in Table 1, produced through the Fragment step described in Section 3.1. In Table 1 high-performing MOFs with three parsed linkers were selected, and their unique linkers parsed by using the MMPA (Matched Molecular Pairs Algorithm) algorithm.
In Table 1, the number of output molecular fragment conformers (last column) is much less than the number of unique linker SMILES (second to last column) for two reasons. First, around 56% of linkers did not pass the valency check, which may be due to the intrinsic limitations in the parsing of MOF linkers.
Second, around 90% of linkers that passed the valency check share similar molecular fragments, and thus many duplicated molecular fragments exist. The successfully parsed molecular fragment conformers are subsequently used for linker generation.
Figure 3: Depictions of the most frequent node-topology pairs in hMOF (left panel) and cumulative distribution functions of the 0.1 bar CO2 capacities (right panel).
Table 1: Data on the hMOF with most frequent node-topology pairs. PW, TM, and HP-MOF stand for paddlewheel, tetramer, and high-performing MOF, respectively.
three linkers SMILES fragment conformers
Cu PW-pcu 29,714 1,458 1,016 3,330 180
Zn PW-pcu 28,529 1,314 834 3,221 162
Zn TM-pcu 43,874 2,129 1,388 3,265 198
Figure 4: Empirical cumulative distribution functions of the 0.1 bar CO2 capacities of hMOF structures at different catenation levels. The x axis is capped at 4 mmol/g to preserve details.
Figure 4 shows the empirical cumulative distribution functions of the CO2 capacities of hMOF structures at different catenation levels. Therein, we observe that a higher percentage of catenated MOFs (cat1, cat2 and cat3) are high-performing, as compared to the uncatenated MOFs (cat0). This result confirms that catenation is an important factor when designing new MOF structures with high CO2 capacity. In Table 5, we summarize the amount of origin shift for each of the four levels of catenation.
Figure 5 shows the pairwise relationships of the CO2 capacities of hMOF structures at the five pressures. We observe a strong correlation of CO2 capacities between 0.05 bar and 0.1 bar, with a Pearson’s correlation coefficient of 0.86. For other pairs of pressures, the CO2 capacities are only weakly correlated, with a decreasing correlation for larger pressure differences. Moreover, the distributions of CO2 capacities at all pressures exhibit long tails at high value ranges, which indicate that the majority of MOFs are low-performing and high-performing MOFs are uncommon.
Figure 5: Pairplot of the CO2 capacities of hMOF structures at different adsorption pressures. The x- and y-axis are adsorption pressures. Different colors indicate MOFs with different node-topology pairs.
4.2 Linker generation and evaluation
For all three MOF types, we start with 540 molecular fragments extracted from high-performing hMOF structures, and use DiffLinker to generate a large pool of 64,800 linkers with the number of sampled connection atoms ranging from 5 to 10. For each linker, sampling is performed 20 times, as described in Section 3.2. Since these linkers only contain heavy atoms, to generate all-atom linker molecules, we apply openbabel to add hydrogen atoms, which results in 56,257 linkers. Next, dummy atom identification is performed to generate information that enables assembly with metal nodes. A total of 16,162 linkers with generated dummy atoms are generated. These linkers are passed through the element filter, which removes linkers that contain S, Br and I elements, further reducing number to 12,305 linkers. The details of the number of linkers in each step are summarized in Table 7.
@@ 4.2 Linker generation and evaluation
We show in Figure 6 the synthetic accessibility score (SAscore) and synthetic complexity score (SCscore) values, defined in Section 3.2.2, for the remaining linkers. We observe that as the number of sampled atoms increases, both distributions generally shift to the right, with the exception of the SAscore distribution with six sampled atoms, which is to the left of that with five atoms. The general trend indicates that linkers become harder to synthesize as the molecules become bigger. This result is expected because as more atoms are sampled, more complex substructures may be present. We note that no linker has zero or very large SAscore or SCscore, values that would indicate unsynthesizability.
@@
Figure 6: Distributions of synthetic accessibility score (SAscore, top row) and synthetic complexity score (SCscore, bottom row) of DiffLinker-generated MOF linkers with dummy atoms. The subfigure titles are node-topology pairs, and the colors listed in the legend indicate the number of sampled atoms, from five to 10 inclusive.
@@
We show in Table 2 the validity, uniqueness, and internal diversity metrics, defined in Section 3.2.2, grouped by number of sampled atoms and node-topology pairs of the corresponding MOFs. The validity column confirms that all generated linkers are valid. The last three columns, when reviewed from top to bottom, reveal a considerable increase in linker uniqueness and a more modest increase in internal diversity as the number of sampled atoms increases. High uniqueness values indicate that our model is generating non-duplicate molecules, whereas high internal diversity values indicate that the generated molecules are chemically diverse. Internal diversity (which indicates how similar/dissimilar a specific linker is to the rest of the population) is computed using two internal diversity scores: IntDiv1 and IntDiv2 [48], also shown in Table 2—see C for details on how these scores are computed. The increase in these metrics as more atoms are sampled is expected because the degrees of freedom of atomic species and their spatial coordinates increase with molecular size.
@@
Table 2: Statistics of the generated linkers that correspond to different MOF types (first column), number of sampled atoms (${\cal{N}}$ in the second column), number of carboxyl linkers (${\cal{C}}$, fourth column), and number of heterocyclic linkers (${\cal{H}}$, fifth column). In total, there are 12,305 linkers with identified dummy atoms.
@@
0.766
@@
Cu paddlewheel-pcu 7 499 383 116 1
0.958
0.836 0.820
Zn paddlewheel-pcu 7 453 329 124 1
0.929
0.829 0.812
Zn tetramer-pcu 7 709 709 0 1
0.958
0.782 0.770
Cu paddlewheel-pcu 8 330 245 85 1
0.992
0.835 0.822
Zn paddlewheel-pcu 8 338 247 91 1
0.978
0.831 0.817
Zn tetramer-pcu 8 550 550 0 1
0.983
0.791 0.780

### v2

@@ Methods
    Fragment: We fragment the linkers produced in Extract step to obtain their chemically relevant fragment-connection atom pairs, which we assemble into a molecular fragment dataset.
The extraction of linkers in Extract step is straightforward because the MOFid of each hMOF structure specifies the SMILES strings of its constituent metal nodes, linkers, as well as a format signature and a topology code [52]. Together, these elements uniquely define the topology and building blocks of a given MOF structure.
Fragment step use MMPA [53] as implemented in DeLinker [54] to generate molecular fragments of a given molecule by breaking chemical bonds between atom pairs. We set the minimum number of connection atoms, the minimum fragment size, and the minimum path length to 3, 5, and 2, respectively. Moreover, we only consider the case where the fragments are at least two atoms away from each other. The chemically relevant fragment-connection atoms pairs are then used to form the molecular fragment dataset.
Figure 7: First component of the GHP-MOFassemble framework. Schematic representation of the Decompose component, which consists of three steps, which we showcase for Zn tetramer-pcu MOFs. First, the Select step selects the Zn tetramer-pcu MOFs in high-performing hMOF structures. Next, the Extract step extracts unique linkers from those MOFs. Finally, the Fragment step generates unique linker fragments. The color scheme of elements is: carbon in grey, oxygen in red, nitrogen in blue, and hydrogen in white. See Table 1 for unique linker fragments of Zn tetramer node.
Generating new MOF structures
@@
We apply the pre-trained diffusion model DiffLinker [31] to generate new linkers based on the molecular fragments outputted by the Decompose component. This model connects the molecular fragments supplied as input (also known as context) with a specified number of sampled atoms. With less sampled atoms, straight chains or branched chains may be obtained, while with more sampled atoms, ring structures may be present. Moreover, during the denoising process, the species of the sampled atoms may change. In this work, we vary the number of sampled atoms from 5 to 10 to ensure that the generated linkers have a diverse set of chemistry and substructures.
@@
DiffLinker applies a denoising process to determine the atomic species and Cartesian coordinates of the sampled atoms by using a decoder neural network architecture named E(3)-Equivariant Graph Neural Network (EGNN) [55]. This graph neural network predicts zero-mean centered Cartesian coordinates noise and one-hot encoding noise of atomic species, and accounts for equivariance due to translation, rotation, and reflection of molecules, i.e., Euclidean Group 3 (E(3)). In this work, we use the pre-trained DiffLinker model that was trained on the GEOM dataset [56], which contains 37 million molecular conformations for over 450,000 molecules. Such diverse training data enables DiffLinker to sample linker molecules with high chemical and structural diversity.
Figure 8: Second component of the GHP-MOFassemble framework. The Generate component of the GHP-MOFassemble framework involves Diffuse and Denoise, Screen and Evaluate, and Assemble steps, as shown here for Zn tetramer-pcu MOFs (see Table 1 for unique linker fragments of Zn tetramer node). First, we generate new linkers via the Diffuse and Denoise step. We then add hydrogen atoms to ensure correct valency of the generated linkers. Next, we identify dummy atoms by replacing the two carbon atoms in the carboxyl groups of each generated linker with dummy atoms. Linkers with dummy atoms then undergo the Screen and Evaluate step, where we remove those with S, Br, and I elements, since these three elements are present in the GEOM dataset but not in the hMOF dataset. As a result, this step reduces the number of potential linkers to 5,162. The generated linkers’ molecular statistics are quantified using five metrics, including SAscore, SCscore, validity, uniqueness, and internal diversity. Finally, in the Assemble step, we build 40,000 new MOFs with Zn tetramer node in pcu topology. As before, the color scheme of elements is: carbon in grey, oxygen in red, nitrogen in blue, and hydrogen in white. See last column of Table 2 regarding this component.
@@
The outputs of DiffLinker are the 3D coordinates and the atomic species of heavy atoms in the linker molecules, rather than SMILES strings or 2D graphs. Since hydrogen atoms are implicit in the DiffLinker model, their 3D coordinates are not outputted by the model. To generate the spatial configurations of all-atom molecules, we employ openbabel to convert the DiffLinker generated molecules to SMILES strings, which in turn are used to generate 3D configurations of linkers with explicit hydrogen atoms. Openbabel uses distance-based heuristics to determine bond connectivity in a given molecule [57], which is critical to the assignment of the number of hydrogen atoms. Finally, after the hydrogen addition step, we identify the dummy atoms which contains information about how the linkers connect with metal nodes.
@@
The diffusion process involves two consecutive steps. The first step adds Gaussian noise to the original data (i.e., x), yielding noisy input (i.e., $z_{t}$ at time t). Next, the denoising step applies neural network-based noise removal operation. Mathematically, the following Markovian properties and equations are satisfied:
@@
Screen and Evaluate
To ensure consistency of elements in hMOF linkers and AI-generated MOF linkers, we manually filter out generated linkers that contain elements not present in the hMOF dataset. Since the pre-trained DiffLinker model we used in this work was trained on the GEOM dataset [56], a total of nine heavy elements may be sampled, including C, N, O, F, P, S, Cl, Br, and I. Among these elements, S, Br, and I do not appear in the hMOF dataset, therefore we remove generated linkers that contain these three elements. For each molecular fragment, we then perform sampling 20 times. Each time the model samples a different molecule may be obtained because of the random and probabilistic nature of denoising process. The probabilistic nature of diffusion model enables it to generate linkers from an extensive linker design space beyond that of the hMOF dataset.
@@
The SAscore, as defined by Ertl and Schuffenhauer [60], is based on analysis of one million PubChem molecules, and combines fragment contributions from molecule substructures with a complexity penalty that accounts for molecular size and for structural features of molecules such as the presence of rings. The SCscore is computed by using a neural network trained on 12 million chemical reactions from the Reaxys database to estimate the number of reaction steps required to produce a molecule [59]. For both SAscore and SCscore, the higher the values are, the more difficult it is to synthesize the linker, hence less desirable. Specifically, the SAscore [60] is defined as the difference of fragment score and complexity penalty:
$\mathrm{SAscore}=\mathrm{fragmentScore}-\mathrm{complexityPenalty},$
where fragment score is calculated by averaging over the contributions of non-zero elements of the Morgan fingerprint for all molecular fragments. The complexity score is calculated based on five components:
$\mathrm{complexityPenalty=-sizePenalty-stereoPenalty-spiroPenalty-bridgePenalty-macrocyclePenalty}.$
The size penalty increases with the number of atoms:
$\mathrm{sizePenalty=nAtoms^{1.005}-nAtoms},$
The stereo penalty is calculated based on the number of chiral centers:
$\mathrm{stereoPenalty=log(nChiralCenters+1)},$
The spiro penalty is calculated based on the number of spiro centers, or atoms that connect two rings together:
$\mathrm{spiroPenalty=\log(nSpiro+1)},$
The bridge penalty term is calculated based on the number of bridgehead atoms, or atoms that belong to two or more rings:
$\mathrm{bridgePenalty=\log(nBridgeheads+1)},$
Lastly, if macrocycles are present, the macroPenalty is:
[…]
otherwise, the macroPenalty is zero. The SCscore [60] is calculated by using a neural network trained on 12 million reactions from the Reaxys database. The synthetic complexity of each molecule is rated from 1 to 5, with 1 being the easiest and 5 the most difficult to synthesize. To evaluate the capability of GHP-MOFassemble to generate a valid, novel, and diverse set of linkers, we leverage the MOSES framework [61] to compute three additional metrics: the fraction of valid linker SMILES strings (validity), the fraction of unique linker SMILES strings (uniqueness), and the dissimilarity of linkers (internal diversity). Validity measures whether atoms in generated molecules have the correct valency, whereas uniqueness quantifies the percentage of any molecule that is different from the rest of molecules.
@@
We show in Figure 9 the synthetic accessibility score (SAscore) and synthetic complexity score (SCscore) values for the remaining linkers. We observe that as the number of sampled atoms increases, both distributions generally shift to the right, with the exception of the SAscore distribution with six sampled atoms, which is to the left of that with five atoms. This general trend indicates that linkers become harder to synthesize as the molecules become bigger. This result is expected because as more atoms are sampled, more complex substructures may be present. We note that no linker has zero or very large SAscore or SCscore, values that would indicate unsynthesizability.
@@
Figure 9: Synthesizability of DiffLinker-generated MOF linkers with dummy atoms. a–c Distributions of synthetic accessibility score (SAscore); and d–f synthetic complexity score (SCscore) of DiffLinker-generated MOF linkers with dummy atoms. We indicate both node-topology pairs, and the number of sampled atoms, from 5 to 10 inclusive.
@@
We show in Table 4 the validity, uniqueness, and internal diversity metrics, grouped by the number of sampled atoms and node-topology pairs of the corresponding MOFs. The validity column confirms that all generated linkers are valid. The last three columns, when reviewed from top to bottom, reveal a considerable increase in linker uniqueness and a more modest increase in internal diversity as the number of sampled atoms increases. High uniqueness values indicate that our model is generating non-duplicate molecules, whereas high internal diversity values indicate that the generated molecules are chemically diverse. Internal diversity (which indicates how dissimilar a specific linker is to the rest of the population) is computed using two internal diversity scores: IntDiv1 and IntDiv2 [61], also shown in Table 4. The increase in these metrics as more atoms are sampled is expected because the degrees of freedom of atomic species and their spatial coordinates increase with molecular size.
@@
Table 4: Statistical properties of AI-generated linkers. Statistics of AI-generated linkers that correspond to different MOF types (first column), number of sampled atoms (${\cal{N}}$ in the second column), number of carboxyl linkers (${\cal{C}}$, fourth column), and number of heterocyclic linkers (${\cal{H}}$, fifth column). In total, there are 12,305 linkers (see Table 2’s Total column, Element filter row) with identified dummy atoms.
@@
0.512
0.693 0.672
Cu paddlewheel-pcu 7 499 383 116 1
0.919
0.751 0.734
Zn paddlewheel-pcu 7 453 329 124 1
0.933
0.758 0.742
Zn tetramer-pcu 7 709 709 0 1
0.866
0.744 0.729
Cu paddlewheel-pcu 8 330 245 85 1
0.947
0.759 0.742
Zn paddlewheel-pcu 8 338 247 91 1
0.943
0.763 0.745
Zn tetramer-pcu 8 550 550 0 1
0.905